# Predicting liquid properties and behavior via droplet pinch-off and machine learning

### v1

@@ Abstract
Figure 1: a) Examples of breaking up droplets at various conditions of Reynolds numbers for different liquids. As expected, pure water shows self-similarity near the pinch-off point and pure glycerin presents long filaments that break-up from the middle. Snapshots of pure glycerin are not to scale. Data domains in terms of Reynolds and Ohnesorge numbers. b) Parametric space of fluid properties and flow characteristics in terms of liquid type and formulation. c) Parametric space expressed in dripping length to nozzle-diameter ratio; many different conditions collapse onto the same $l/D.$
Machine learning (ML) is a branch of computer modeling that learns patterns from large amounts of data and makes predictions, or decisions, without prior knowledge beyond what it has learned from this data. In recent years, ML has been applied across various physics topics, including fluid dynamics(35, 36, 37), droplet physics (38, 39, 40, 41, 42), soft matter (43, 44),and materials science (45). These studies highlight the potential of ML as a tool to complement traditional experimental and theoretical approaches in science. Specifically, supervised learning trains models on labelled datasets (data that has been manually classified or characterised), mapping inputs to known outputs, which is particularly effective for prediction tasks. In contrast, unsupervised learning operates on unlabelled data to identify hidden patterns or underlying structures, i.e. clusters. Both strategies offer valuable means of extracting knowledge from complex experimental observations. In addition, recent advances in computer vision(46) provide powerful tools for analysing image-based experimental data (47), which are highly relevant for the study of droplet dynamics and pinch-off processes where high-speed imaging is often used in conjunction with image analysis to extract information from images(19).
ML methods have been applied to various systems involving droplet formation, examples of these are on direct inkjet writing(48), droplet impact, spreading and bouncing on nanostructure substrates(49), spreading on various materials (50), and droplet splashing (51). Considering first the forward-prediction paradigm, in which ML models predict droplet behaviour from known material and operating conditions, Kim et al.(52) trained supervised regressors to predict the jetting behaviour of viscoelastic inks from rheological and drive parameters, with later work (53) using reinforcement learning to optimise actuation waveforms, dynamically improving jet stability without an explicit physical model. Additionally, Chen et al.(39) employed supervised learning to predict the optimum printable biomaterial formulations for direct ink writing, and Li et al.(41) developed a multi-parameter droplet optimisation model including flow and material parameters. Turning to image-based studies, Kratz and Kierfeld (40)introduced an image-based ML approach for pendant-drop tensiometry, demonstrating that the contour of a static droplet can be used to infer surface tension with the method restricted to steady pendant configurations. Pierzyna et al. (51) developed a data-driven model for the splashing thresholds and Au-Yeung et al. (49) developed a method to predict the maximum spreading and impact morphology of droplets on nanostructured surfaces.
In summary, previous works have predominantly used ML to control, or classify, droplet behaviour given known physical inputs. In contrast, the present work uses supervised learning to tackle the inverse problem of inferring fluid properties directly from droplet morphology, in addition we use unsupervised learning to gain insight into the underlying physics of this process.
Results and Discussion
Experimental Observations
Figure 1a presents examples of snapshots captured at the frame immediately prior to breakup. The images are organised by nozzle diameter $D$ and the liquid composition. As can be seen, water (a near inviscid fluid) displays a characteristic sharp neck that previous works have confirmed takes a 18 degree angle. In contrast, the 50% mixture forms a longer, more gradually evolving neck, which indicates viscosity is playing a role in the dynamics. Interestingly, regardless of having very similar Reynolds numbers, pure water and 50% glycerin (with $D=0.25$ mm) have visibly different shapes, with the glycerin solution having a larger bulb. Pure glycerin shows a slender, quasi–cylindrical long filament, that pinches off at the middle, consistent with a viscous–dominated dynamics. The Reynolds numbers listed beneath the panels increase as $D$ decreases, e.g. for water $Re=0.77\rightarrow 2.56\rightarrow 6.79$, and for glycerin $Re=0.001\rightarrow 0.023\rightarrow 0.080$). As can be seen, increasing viscosity (decreasing $Re$) shifts the minimum neck radius upstream and increases the break-up length to nozzle diameter aspect ratio.
The parameter map in Figure 1b identifies different liquids within the ReReOhOh plane. As expected, methanol is found at the lowest $Oh$ values due to its low viscosity. In contrast, aqueous ethanols are found in the low–$Oh$, and moderate to high $Re$, region due to their low viscosities and surface tensions. Glycerin–water mixtures occupy intermediate $Oh$ values across a wide range of $Re$, with pure glycerin, and high–viscosity silicone oils, situating at high $Oh$ and low $Re$. Our accessible domain is limited in the low left by very low viscosity and very slow flow rates, and in the upper right by the maximum pumping pressure, which limits the flow rates. We argue that our domain of experimental conditions is sufficiently extensive as it includes very low viscosity liquids (methanol and water), intermediate viscosities (solutions) and very high viscous fluids (silicon oil and glycerin), and very low surface tension (silicon oils) and high surface tension liquids (water).
Figure 1c provides an alternative perspective by color-mapping the break-up length-to-nozzle diameter ratio $L/D$ across the ReReOhOh space. As noted in the Introduction, the $L/D$ ratio has been used to determine whether a liquid filament will break up or not, and to distinguish dripping from jetting conditions (33, 34). From our results (the colour redundancy in Figure 1c), various $Re$ and $Oh$ conditions lead to the same dimensionless break-up length; it is hence clear that the aspect ratio alone is insufficient to predict liquid properties, or a flow characteristic.
@@
Figure 2: Predicted vs. true values on the test set for Models 1–3 (MLP). Model 1 (a) predicts the viscosity, Model 2 (b) predicts surface tension, and Model 3 predicts both the viscosity (c) and the surface tension (d).
@@
Figure 3: Predicted vs. true values on the test set for Models 1–3 (XGBoost).Model 1 (a) predicts the viscosity, Model 2 (b) predicts surface tension, and Model 3 predicts both the viscosity (c) and the surface tension (d).
Supervised learning results
@@
Model 1, implemented as a multilayer perceptron(MLP), predicts the viscosity from the other liquid properties, the flow characteristics, and the image-derived shape features. Figure 2a uses a logarithmic scale and covers more than three orders of magnitude; we obtain $R^{2}=0.9428$. In contrast, with the same inputs, gradient boosting decision trees(XGBoost) reaches $R^{2}=0.9978$ (Fig. 3a), confirming that the viscosity signal is extremely well aligned with the latent geometry. We note XGBoost is known to handle well small datasets.
@@
Model 3 is the multitask setting, predicting both viscosity and surface tension together. Our MLP yields $R^{2}_{\mu}=0.9361$ for viscosity and $R^{2}_{\sigma}=0.9854$ for surface tension. The XGBoost version of the same task improves the numbers to $R^{2}_{\mu}=0.9831$ and $R^{2}_{\sigma}=0.9995$ (Fig. 3c and 3d). Despite having to predict two parameters simultaneously, the multi-task setting maintains essentially the same fidelity as the single-task models.
Model 4 differs from the others in that it predicts the shape (the latent vector) directly from physical parameters, without any image as input. Using the MLP, we previously obtained $R^{2}=0.9145$ on the test set when comparing predicted latent vector. The XGBoost model reaches $R^{2}=0.9713$ in the normalized latent space. Because the decoder is a nonlinear map from latent coordinates to pixels, part of this advantage is smoothed out after decoding, but the high latent-space $R^{2}$ confirms that a large fraction of the image variability can be explained directly from the physical descriptors. Representative predictions are shown in Fig. 4a and Fig. 4b , we can see MLP and XGBoost models obtained good prediction. Both models reproduce the global morphology reliably: the bulb radius, neck position, and overall aspect ratio match the ground truth. From the differences in the figures, it can be seen that MLP better maintained the continuity of the lines, but slightly underestimated the terminal bulbs, while XGBoost underestimated the thickness of the filament neck.
We noted that, across all four supervised tasks, both the MLPs and the XGBoost models consistently achieve test-set $R^{2}>0.9$, and in the viscosity case the boosted model is effectively at the noise floor ($R^{2}\approx 0.998$). This compares favourably with prior work (38, 42). Remarkably, the tree-based models (XGBoost) match or even slightly outperform the MLPs on several tasks. These numbers show that once a good latent vector is available, even a relatively shallow ensemble of boosted trees can exploit it extremely well.
Refer to captiona)
Refer to captionb)
Refer to captionc)
@@
Figure 4: Model 4 predictions: experimental vs. predicted droplet shapes for both (a) MLP and (b) XGBoost variants. (c) Unsupervised learning results: average droplet shapes in each GMM and K-Means cluster.
Unsupervised learning results

### v2

@@ III.2 Image Analysis
Figure 3: Data domains in terms of the Reynolds and Ohnesorge numbers. a) Parametric space of fluid properties and flow characteristics in terms of the liquid type and formulation; most liquids correspond to aqueous solutions of ethanol and glycerin. b) Parametric space in terms of the dripping length to nozzle inner diameter ratio (I/D); as seen, the ratio is not a simple function of $Oh$ and ReRe as many conditions lead to the same aspect ratio.
a) Refer to caption b) Refer to caption
Figure 4: Autoencoder error analysis : (a) Relative reconstruction error $\|D_{\phi}-E_{\phi}\|/\|E_{\phi}\|$ on the validation set versus the latent dimension $d$ of the shape vector ($z\in\mathbb{R}^{d}$). As seen, the minimum error is found at $d=14$, which is the value used in our models. (b) Qualitative comparison of a representative boundary: ground truth (original) vs. autoencoder output (reconstructed). Here, $Re=7.943$ and Oh=0.004Oh=0.004(90% water + 10% glycerin).
IV Machine learning methodology
A common foundation across both approaches is the use of a Convolutional Neural Network Autoencoder (CNN-AE) [46] to extract low-dimensional representations of the droplet snapshots, in the form of a latent shape vector z. The structure of the autoencoder can be found in Table 2 in Appendix A. The primary motivation for employing a CNN is its ability to filter and de-noise raw image data while reducing the input dimension. These features are then used as model inputs in the supervised learning tasks, or as clustering coordinates in the unsupervised learning analyses. Our dataset is constructed from the droplet boundary-only images previously described. We note the experimental images were obtained as high-resolution monochromes, but they were neither square nor of the resolution required. Consequently, to ensure consistent input sizes for the autoencoder, we padded the original images with a white background to transform them into square images of a $875\times 875$ ($x\in\mathbb{R}^{875\times 875}$) size; this padding preserves the original resolution and aspect ratio of the droplet shapes without any downsampling or interpolation. CNNs require that all images within a batch share the same spatial dimensions; padding was therefore applied to obtain square inputs, standardizing image size and avoiding distortions caused by rescaling.
From our dataset of 840 samples we randomly selected 600 samples for training, 120 samples for validation, and 120 samples for the testing the autoencoder and the other methods. At each step, these groups of samples remained isolated from each other to avoid data leakage. The network was trained using the Adam optimizer [56] with a learning rate of $2.5\times 10^{-4}$. An early stopping criterion was applied: if the validation loss did not decrease for 30 consecutive epochs, training was terminated. Under these conditions, the training converged after 140 epochs with a batch size of 32. Loss curves for the autoencoder are provided in Fig.11 in Appendix A.
Figure 5: Supervised learning framework (XGBoost). Model inputs consist of physical parameters and, when required, the 14-dimensional latent droplet representation. The models predict viscosity (Model 1), surface tension (Model 2), viscosity and surface tension (Model 3), or the 14-dimensional shape descriptor (Model 4).
It should be noted that the data used for validation are not the data used for final, out of sample testing; the test loss is only evaluated once training is completed and provides an unbiased measure of the final model performance. A learning rate scheduler reduces the learning rate by a factor of 0.5 after a quarter of the training epochs. To improve generalization, random Gaussian noise with a standard deviation of 0.02 is added to the input images during training. The model uses binary cross-entropy (BCE) loss, as the output is normalized. Mathematically, the autoencoder consists of an encoder $E_{\theta}$ and decoder $D_{\phi}$, parameterized by $\theta$ and $\phi$, respectively. Given an input image $x$, the BCE reconstruction loss is defined as
$\begin{split}\mathcal{L}_{\text{BCE}}(x,D_{\phi}(E_{\theta}(x)))&=-\frac{1}{N}\sum_{i=1}^{N}\Big[x_{i}\log(D_{\phi}(E_{\theta}(x))_{i})\\ &\quad+(1-x_{i})\log(1-D_{\phi}(E_{\theta}(x))_{i})\Big].\end{split}$
where $x$ is the ground truth image, $D_{\phi}(E_{\theta}(x))$ is the reconstructed image, both with pixel values in $[0,1]$, and $N$ is the total number of pixels. This loss encourages the model to accurately predict pixel intensities in the reconstructed image [57]. The relative reconstruction error was computed as
$\text{RE}(x,D_{\phi}(E_{\theta}(x)))=\frac{\lVert x-D_{\phi}(E_{\theta}(x))\rVert_{2}}{\lVert x\rVert_{2}}.$
The validation set was then used to evaluate the reconstruction performance of the autoencoder. Figure 4a shows the relative reconstruction error as a function of the latent dimension. Our results indicate that the relative error decreases rapidly as the latent dimension $d$ (the length of vector z) increases, and plateaus around $d=14$, indicating that adding more dimensions does not improve accuracy. Consequently, in our methods, we set the autoencoder to produce a vector z of size $d=14$. We note that we have found a low-dimensional representation with a dimension in $\mathcal{O}(10^{1})$, starting from an input dimension of $\mathcal{O}(10^{5})$; thus a reduction of four orders of magnitude is achieved. A representative reconstruction example is found in Fig. 4b, contrasting the original contour and the reconstruction (with $d=14$), and demonstrating that the autoencoder effectively captures the overall droplet geometry.
IV.1 Predicting liquid properties & droplet shape with supervised learning
In this section, we describe the supervised models used to infer either one or two fluid properties (viscosity and/or surface tension) from images, or to predict the droplet shape from a set of known physical parameters. All supervised learning models were implemented using two types of algorithms: multilayer perceptron (MLP) and gradient boosting decision trees (XGBoost) [58], which differ in their learning paradigms. The four XGBoost tasks, or models, considered are summarized in Figure 5, and details the MLP model are found in in Appendix B. Neural networks learn smooth continuous mappings, while XGBoost captures nonlinear feature interactions through ensembles of decision trees—allowing a complementary comparison between deep and tree-based methods. Models 1–3 combine the autoencoder latent vector z with physical parameters (nozzle diameter, flow rate, density, viscosity, and surface tension) to predict fluid properties. Model 4 performs the inverse mapping, predicting droplet shape from the same physical parameters, thereby establishing a bidirectional link between geometry and material behaviour. In addition, the same train, validation, and test split as used for the autoencoder was used here, with the same strict separation of these sets, in order to prevent data leakage. Each model was trained using the Adam optimizer with an initial learning rate of $10^{-3}$ and the mean-squared-error (MSE) loss function,
$\mathcal{L}_{\mathrm{MSE}}=\frac{1}{N}\sum_{i=1}^{N}\|\hat{y}_{i}-y_{i}\|_{2}^{2}.$
where $N$ denotes the number of samples in a batch. For Models 1–3, the output values were log-transformed during training to improve numerical stability. A batch size of 32 was employed, and a StepLR scheduler was applied with a step size of 100 epochs and a decay factor of $0.8$. Early stopping with a patience of 100 epochs was applied based on validation loss, and the model weights corresponding to the lowest validation loss were restored for testing. For targets with large dynamic range (e.g., viscosity), the outputs were log-transformed during training to improve numerical stability, and then mapped back to the physical scale for evaluation.
Models 1 and 2 are formulated as single-task regressors. In Model 1, the inputs consist of the droplet shape latent vector obtained from the CNN encoder, together with the density, diameter, flow rate and the surface tension; the network is trained to predict the viscosity $\mu$. Model 2 receives the same latent vector, plus density, diameter, flow rate, and the viscosity $\mu$, to predict the surface tension $\sigma$. These two models served as baselines, inferring one unknown physical property of the example given the droplet shape (via the encoder representation) and the known other physical properties. Model 3 extends this approach to a multi-task setting. Here, the inputs are only the latent shape vector and the control parameters of density, nozzle size and flow rate, and the network simultaneously outputs both viscosity $\mu$ and surface tension $\sigma$.This design allows the model to learn a shared network structure for the two correlated fluid properties, and provides a more realistic scenario in which neither property is known a priori. Together, Models 1 to 3 provide a progressive comparison between single-task and multi-task learning for fluid property prediction. The performance of the models was assessed by the coefficient of determination $R^{2}$, defined as
$R^{2}=1-\frac{\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}},$
where $y_{i}$ denotes the true (experimentally obtained) values, $\hat{y}_{i}$ the (ML-obtained) predicted values, and $\bar{y}$ the mean of the true values.
Figure 6: Examples of breaking up droplets at various conditions of Reynolds numbers for different liquids. As expected, pure water shows self-similarity near the pinch-off point and pure glycerin presents long filaments that break-up from the middle. Snapshots of pure glycerin are not to scale.
IV.2 Uncovering natural groupings in droplet behaviour with unsupervised learning
In addition, we have employed unsupervised learning to directly uncover naturally occurring group behaviour in droplet dynamics from the observed shapes and imposed flow conditions, without requiring prior knowledge of the underlying fluid properties.
This work is motivated by experimental scenarios in which viscosity or surface tension are not readily measurable, yet droplet morphology and control parameters are accessible, e.g. in systems where liquid volumes are too small. Specifically, we extracted low-dimensional latent features $z\in\mathbb{R}^{14}$ from the encoder, with the latent dimension $d=14$ adopted from Section IV, where it was identified as the optimal representation size. These latent features were combined with two directly controlled experimental parameters, nozzle inner diameter and flow rate, for clustering. Notably, we deliberately excluded any information related to the fluid properties, as these quantities are often difficult to measure in practice. We applied both K-Means and Gaussian Mixture Models (GMM) with cluster numbers $K$. To identify optimal settings, we systematically tested $K\in\{2,7\}$. The choice of the upper bound $K=7$ reflects a practical consideration: while larger values of $K$ might further partition the data, excessively fine-grained clustering may produce groups which lack clear physical interpretation in the context of droplet dynamics. Thus, restricting $K$ to moderate values ensures a balance between capturing heterogeneity in the dataset and maintaining physically meaningful cluster assignments.
In order to assess clustering quality, we employed two complementary metrics, the Silhouette Score [59] and the Davies-Bouldin Index (DBI) [60]. The Silhouette Score quantifies how well a sample fits within its assigned cluster compared to the nearest alternative cluster. For a data point $z_{i}$, the score is defined as
$s(z_{i})=\frac{b(z_{i})-a(z_{i})}{\max\{a(z_{i}),\,b(z_{i})\}},$
$DBI=\frac{1}{K}\sum_{i=1}^{K}\max_{j\neq i}\left(\frac{s_{i}+s_{j}}{d_{ij}}\right),$
where $s_{i}$ is the average distance between points in cluster $i$ and its centroid, and $d_{ij}$ is the distance between the centroids of clusters $i$ and $j$. Lower DBI values indicate better clustering performance, favoring compact and well-separated clusters. The algorithm settings for GMM and K-Means are provided in Appendix C, and the evaluation of the cluster number $K$ is discussed in Appendix D.
@@ IV.2 Uncovering natural groupings in droplet behaviour with unsupervised learning
Figure 7: Predicted vs. true values on the test set for Models 1–3 (MLP). Model 1 (a) predicts viscosity, Model 2 (b) predicts surface tension, and Model 3 predicts both viscosity (c) and surface tension (d).
V Experimental observations
Figure 6 presents examples of snapshots captured at the frame immediately prior to breakup. The images are organised by nozzle diameter $D$ and the liquid composition. As can be seen, water (a near inviscid fluid) displays a characteristic sharp neck that previous works have confirmed takes a 18 degree angle. In contrast, the 50% mixture forms a longer, more gradually evolving neck, which indicates viscosity is playing a role in the dynamics. Interestingly, regardless of having very similar Reynolds numbers, pure water and 50% glycerin (with $D=0.25$ mm) have visibly different shapes, with the glycerin solution having a larger bulb. Pure glycerin shows a slender, quasi–cylindrical long filament, that pinches off at the middle, consistent with a viscous–dominated dynamics. The Reynolds numbers listed beneath the panels increase as $D$ decreases, e.g. for water $Re=0.77\rightarrow 2.56\rightarrow 6.79$, and for glycerin $Re=0.001\rightarrow 0.023\rightarrow 0.080$). As can be seen, increasing viscosity (decreasing $Re$) shifts the minimum neck radius upstream and increases the break-up length to nozzle diameter aspect ratio.
The parameter map in Figure 3a identifies different liquids within the ReReOhOh plane. As expected, methanol is found at the lowest $Oh$ values due to its low viscosity. In contrast, aqueous ethanols are found in the low–$Oh$, and moderate to high $Re$, region due to their low viscosities and surface tensions. Glycerin–water mixtures occupy intermediate $Oh$ values across a wide range of $Re$, with pure glycerin, and high–viscosity silicone oils, situating at high $Oh$ and low $Re$. The accessible domain of our setup is limited in the low left by very low viscosity and very slow flow rates, and in the upper right by the maximum pumping pressure, which limits the flow rates. We argue that our domain of experimental conditions is sufficiently extensive as it includes very low viscosity liquids (methanol and water), intermediate viscosities (solutions) and very high viscous fluids (silicon oil and glycerin), and very low surface tension (silicon oils) and high surface tension liquids (water).
Figure 3b provides an alternative perspective by color-mapping the break-up length-to-nozzle diameter ratio $L/D$ across the ReReOhOh space. As noted in the Introduction, the $L/D$ ratio has been used to determine whether a liquid filament will break up or not, and to distinguish dripping from jetting conditions [33, 34]. From our results (the colour redundancy in Figure 3b), various $Re$ and $Oh$ conditions lead to the same dimensionless break-up length; it is hence clear that the aspect ratio alone is insufficient to predict liquid properties, or a flow characteristic.
@@ V Experimental observations
Figure 8: Predicted vs. true values on the test set for Models 1–3 (XGBoost). Model 1 (a) predicts viscosity, Model 2 (b) predicts surface tension, and Model 3 predicts both viscosity (c) and surface tension (d).
VI Machine learning results
VI.1 Supervised learning results
@@
Model 1, implemented as a multilayer perceptron (MLP), predicts the viscosity from the other liquid properties, the flow characteristics, and the image-derived shape features. Figure 7a uses a logarithmic scale and covers more than three orders of magnitude; we obtain $R^{2}=0.9428$. In contrast, with the same inputs, gradient boosting decision trees (XGBoost) reaches $R^{2}=0.9978$ (Fig. 8a), confirming that the viscosity signal is extremely well aligned with the latent geometry. We note that XGBoost is known to handle well small datasets.
@@
Model 3 is the multitask setting, predicting both viscosity and surface tension together. Our MLP yields $R^{2}_{\mu}=0.9361$ for viscosity and $R^{2}_{\sigma}=0.9854$ for surface tension. The XGBoost version of the same task improves the numbers to $R^{2}_{\mu}=0.9831$ and $R^{2}_{\sigma}=0.9995$ (Fig. 8c and 8d). Despite having to predict two parameters simultaneously, the multi-task setting maintains essentially the same fidelity as the single-task models.
Refer to caption(a)
Refer to caption(b)
Refer to caption(c)
@@
Figure 9: Model 4 predictions: experimental vs. predicted droplet shapes for both (a) MLP and (b) XGBoost variants. (c) Unsupervised learning results: average droplet shapes discovered by GMM and $K$-Means clustering in latent space.
Model 4 differs from the others in that it predicts the shape (the latent vector) directly from physical parameters, without any image as input. Using the MLP, we previously obtained $R^{2}=0.9145$ on the test set when comparing predicted latent vector. The XGBoost model reaches $R^{2}=0.9713$ in the normalized latent space. Because the decoder is a nonlinear map from latent coordinates to pixels, part of this advantage is smoothed out after decoding, but the high latent-space $R^{2}$ confirms that a large fraction of the image variability can be explained directly from the physical descriptors. Representative predictions are shown in Fig. 9a and Fig. 9b, we can see that both MLP and XGBoost models provide good prediction. Both models reproduce the global morphology reliably: the bulb radius, neck position, and overall aspect ratio match the ground truth. From the differences in the figures, it can be seen that MLP better maintained the continuity of the lines, but slightly underestimated the terminal bulbs, while XGBoost underestimated the thickness of the filament neck.
We noted that, across all four supervised tasks, both the MLPs and the XGBoost models consistently achieve test-set $R^{2}>0.9$, and in the viscosity case the boosted model is effectively at the noise floor ($R^{2}\approx 0.998$). This compares favourably with prior work [38, 42]. Remarkably, the tree-based models (XGBoost) match or even slightly outperform the MLPs on several tasks. This is likely because the physical properties in our dataset, e.g. viscosity and surface tension, take only discrete values, determined by the limited categories of liquids used. While the MLP attempts to learn a smooth, continuous mapping, XGBoost is naturally better suited for capturing piecewise-constant relationships arising from such discrete labels. The performance of all 4 models show that once a good latent vector is available, even a relatively shallow ensemble of boosted trees can exploit it extremely well.
VI.2 Unsupervised learning results